OpenStudy (anonymous):

Blake knows that one of the solutions to x2 – 6x + 8 = 0 is x = 2. What is the other solution?

OpenStudy (anonymous):

You need to factor this out.

OpenStudy (anonymous):

x^2 - 6x + 8=0 can be factored out into (x-4)(x-2)

OpenStudy (anonymous):

We know that one answer is 2, so what is the other answer?

OpenStudy (gorv):

product of solution = 8

OpenStudy (gorv):

sum of solution=6

OpenStudy (gorv):

you have one solution so 2+a=6 or 2*a=8 solve any of two

OpenStudy (anonymous):

This is what I got

OpenStudy (anonymous):

But not sure if it's right

OpenStudy (anonymous):

That is correct, but we already know 2 is a solution. We need to find the second one. gorv used Vieta's to solve it, which is perfectly fine

OpenStudy (gorv):

what you uploaded ..that is the proof that 2 is solution but we need to to find another